OpenStudy (anonymous):

2m/5 = 1/3 (2m - 12)

OpenStudy (anonymous):

solving for m?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well first you distribute the 1/3

OpenStudy (anonymous):

and that is

OpenStudy (anonymous):

kind of stuck on it

OpenStudy (anonymous):

so 2m/5=(2/3m-4)

OpenStudy (anonymous):

ok then i .......

OpenStudy (anonymous):

then you multiply both sides by 5

OpenStudy (anonymous):

to what

OpenStudy (anonymous):

so 2m=10/3m-20

OpenStudy (anonymous):

to both sides

OpenStudy (anonymous):

because youre getting rid of it in the denominator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then divide 2 from both sides

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

or do i subtract

OpenStudy (anonymous):

you could or you could subtract 10/3m from both sides

OpenStudy (anonymous):

oh wait a sec i think i messed up

OpenStudy (anonymous):

actually you would multiply both sides by 3m to get rid of the denominator

OpenStudy (anonymous):

:S

OpenStudy (anonymous):

jk i have no idea what im doing hold on lemme think about it

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

yeah i have no idea

OpenStudy (anonymous):

ok so :S

OpenStudy (anonymous):

what do i do

OpenStudy (anonymous):

i tried to divide both sides by 1/3 in the beginning by multiplying by the reciprocal and i got 6m/5=2m-12

OpenStudy (anonymous):

but not sure...

OpenStudy (anonymous):

I FIGURED IT OUT

OpenStudy (anonymous):

then you multiply both sides by 5

OpenStudy (anonymous):

so 6m=10m-60

OpenStudy (anonymous):

then you get the right stuff on the right sides

OpenStudy (anonymous):

so -4m=-60

OpenStudy (anonymous):

then divide -60 by -4

OpenStudy (anonymous):

annnnndddddd m=15

OpenStudy (anonymous):

...youre welcome *takes a bow*